b)  Answer the following stoichiometry questions by referring to the equation below:

2       

KClO3               2 KCl + 3 O2

i.        If 1.50 mol of KClO3 decomposes, what is the mass of O2 that will be produced? 

1
Expert's answer
2021-03-15T09:29:11-0400

This is a problem of determining mole ratios in the stoichiometric equations.

(i) Mass of O2

We have moles of KClO3 = 1.50mol

Mole ratios of KClO3 : O2 = 2:3

Thus moles of O2 "=\dfrac{3}{2}x1.50mol"


= 2.25mol of O2


Moles "=\dfrac{Mass}{MM}"


Therefore Mass of O2 = Moles x MM

= 2.25mol x 32g/mol

= 72g

(ii) Moles of KClO3


Moles of O2 "=\dfrac{mass}{MM} = \dfrac{80g}{32g/mol} = 2.5mol"


Mole ratios of O2:KClO3 = 3:2


Moles of KClO3 "=\dfrac{2}{3}x 2.5mol = 1.67 mol"


(ii)Mass of KClO3

We are given moles of KCl = 2.75mol

Mole ratios of KCl : KClO3 = 2:2

Thus moles of KClO3 "=\dfrac{2}{2}x2.75mol = 2.75 mol"


Mass = Moles x MM

= 2.75mol x 122.55g/mol

= 337.013g